# Fragmentation of nuclear remnants in electron-nucleus collisions at high energy

### v2

@@
Abstract: It is widely accepted that the various internal structures of an excited nucleus can lead to different topological configurations of nuclear fragments during fragmentation. However, these internal structures may transition from one configuration to another with varying probabilities. Utilizing a partitioning method based on equal (or unequal) probabilities—without incorporating the alpha-cluster ($\alpha$-cluster) model—allows for the derivation of diverse topological configurations of nuclear fragments resulting from fragmentation. Subsequently, we predict the multiplicity distribution of nuclear fragments for specific excited nuclei, such as 9Be, 12C, and 16O, which can be formed as nuclear remnants in electron-nucleus ($eA$) collisions at high energy. According to the $\alpha$-cluster model, an $\alpha$-cluster structure should take precedence among different internal structures; this may result in deviations in the multiplicity distributions of nuclear fragments with charge $Z=2$, compared to those predicted by the partitioning methods. The findings presented herein could serve as a reference for validating the existence of $\alpha$-cluster structures.
Keywords: $\alpha$-cluster structure; nuclear structure; multiplicity distribution of fragments; the Electron-Ion Collider
@@ 1 Introduction
High-energy nuclear collisions represent a significant area of research within modern physics [2, 3, 4, 5, 6]. In these interactions, numerous particles are predominantly generated within participant regions while multiple fragments are primarily emitted from spectator regions when available. During multi-fragment emission processes, it is anticipated that spectators will form an excited nucleus. This excited nucleus subsequently undergoes fragmentation into various components [7, 8, 9, 10]. Such fragmentation reveals rich internal structures within the nucleus where protons and neutrons can combine appropriately to create intermediate configurations [11, 12, 13, 14]. When two protons and two neutrons coalesce into such an intermediate structure, it is referred to as an alpha-cluster ($\alpha$-cluster) structure [15, 16, 17, 18]. It is possible that two or more $\alpha$-cluster structures are existent in a heavy nucleus [19, 20, 21, 22, 23]. According to the $\alpha$-cluster model [24, 25, 26, 27, 28], the $\alpha$-cluster structure should have much higher probability than other intermediate structures.
@@
In this work, both equal and unequal probability partitioning methods are employed to derive various configurations of nuclear fragments from excited states such as 9Be, 12C, and 16O, which are expected to result from electron-nucleus ($eA$) collisions at the EIC [33]. The multiplicity distributions for all fragments as well as those with charge $Z$ will subsequently be obtained. In particular, the probability of multi-$\alpha$ or multi-He configuration or channel can be obtained, which may serve as the baseline for judging about the $\alpha$-cluster structure.
@@ 2 Various configurations of nuclear fragments
[…]
@@
[…]
@@
Considering the Fermi momentum of a nucleon within the nucleus, which is approximately 0.25 GeV$/c$, and the momentum per nucleon of the incident nucleus being 10 GeV$/c$, decay protons and other fragments are expected to be emitted within a forward cone characterized by a polar angle $\theta_{0}=25$ mrad. This corresponds to a pseudorapidity of $\eta=-\ln\tan(\theta_{0}/2)=4.38$. Furthermore, recoil protons can be distinguished from decay protons since recoil protons participate in multi-scattering processes which result in much larger scattering angles. It is assumed that this emission will span a wide range from nearly 0 (corresponding to $\eta=\infty$) up to approximately $10\theta_{0}$ (which corresponds to $\eta=2.07$). Indeed, it cannot be excluded that some recoil protons may have very small scattering angles, leading to exceptionally large pseudorapidities due to the influence of leading nucleons.
If the two types of protons are assumed to emit isotropically in their respective rest frames, they approximately follow Gaussian $\eta$ distributions with a common standard deviation ($\sigma_{\eta}\approx 0.91$). The decay protons predominantly distribute within the range of $8.024.38<\eta<4.38+4\sigma_{\eta}=$ 8.02, while the recoil protons primarily occupy the range of $5.712.07<\eta<2.07+4\sigma_{\eta}=$ 5.71. It is evident that there exists some overlap between decay and recoil protons in the forward cone. Although most recoil protons may have emission angles exceeding 25 mrad due to multiple scattering processes, we cannot entirely dismiss the possibility of them appearing within the forward cone. To more effectively distinguish between decay and recoil protons, one could study their energies; generally, the energy of a decay proton is nearly equal to that of an incident nucleus per nucleon, whereas the energy of a recoil proton should be lower.
@@
Table 1: The multiplicity, $N_{F}$, of all fragments and the multiplicity, $N_{Z}$, of the fragments with charge $Z$ in various configurations in excited 9Be fragmentation, where only the charge conservation is considered in the fragmentation. In the equal probability partitioning method, the fragment Be is defaulted with 50% probability to be 8Be, and in the unequal probability partitioning method, the fragment Be is defaulted with a given chance $\{M_{2}(2{\rm He})/[M_{2}(2{\rm He})+M_{2}({\rm Be})]=1/3\}$ to be 8Be, where 8Be is unstable and can decay into 2He, which is listed in the bracket, and causes $N_{F}$ to $N_{F}+1$ and $N_{Z=2}$ to $N_{Z=2}+2$. Here, the changeable $N_{F}$ and $N_{Z=2}$ are shown in the table by $+1$ and $+2$, respectively. The probabilities $f_{1}$ and $f_{2}$ of each channel obtained by the equal and unequal partitioning methods are listed respectively.
@@ 3 Multiplicity distributions of nuclear fragments
The normalized multiplicity distributions, $(1/n)(dn/dN_{F})$ or $(1/n)(dn/dN_{Z})$, in excited 9Be, 12C, and 16O fragmentation are displayed in Figures 1, 2, and 3, respectively. The solid (dashed) histograms represent the results from the equal (unequal) probability partitioning method. Figures 1(a), 2(a), and 3(a) are for the multiplicity distributions of all fragments. The multiplicity distributions of the fragments with different $Z$ are shown in different panels, where $Z=1$–4 in Figures 1(b)–1(e), $Z=1$–6 in Figures 2(b)–2(g), and $Z=1$–8 in Figures 3(b)–3(i) are for excited 9Be, 12C, and 16O fragmentation, respectively. It is worth noting that the predicted frequency distributions in Figures 1–3 are precise numerical values (fractions) in each bin under a given scenario.
@@
In this context, we assume that the experimental percentage of multi-He events follows a Gaussian distribution with standard deviation $\sigma$, predominantly concentrated within the range [$0,4\sigma$] and centered around an expected value of $2\sigma$, which serves as our baseline. If the experimental percentage exceeds $4\sigma$, defined as twice the baseline and serving as our threshold for judgment, one can draw conclusions regarding the existence of 3He or $\alpha$ clustering with over 95% confidence level. According to this line, we conclude that 3He or $\alpha$ clustering exists in excited 9Be formed during peripheral collisions between 10B and nuclear emulsion at $E_{\rm beam}=1$ GeV/nucleon.
@@
To ensure an accurate description as possible, the heaviest fragments produced in an event—considered remnants of excited nuclear fragmentation—should be excluded from analysis. For genuine evaporation products, it must be acknowledged that they originate from the fragmentation process involving a smaller excited nucleus; thus, the partitioning methods should be reapplied specifically for this smaller nucleus. Whether a liquid-gas phase transition occurs in the overall or local area, the proportion of light fragments with $Z=1$ should exceed twice the baseline values when applying Poisson distribution to the considered probabilities. Furthermore, if experimental measurements fall within the theoretical uncertainty range, fragmentation may be interpreted as a consequence of a general stochastic process.

### v4

@@
Abstract: Utilizing a partitioning method based on equal (or unequal) probabilities—without incorporating the alpha-cluster ($\alpha$-cluster) model—allows for the derivation of diverse topological configurations of nuclear fragments resulting from fragmentation. Subsequently, we predict the multiplicity distribution of nuclear fragments for specific excited nuclei, such as 9Be, 12C, and 16O, which can be formed as nuclear remnants in electron-nucleus ($eA$) collisions at high energy. Based on the $\alpha$-cluster model, an $\alpha$-cluster structure may result in deviations in the multiplicity distributions of nuclear fragments with charge $Z=2$, compared to those predicted by the partitioning methods. Furthermore, in the framework of Tsallis statistics, the nonextensive generalized temperature, entropy index, and $q$-entropy are obtained from the multiplicity distribution of nuclear fragments with given charge number. Our work shows that fragmentation of nuclear remnants in electron-nucleus collisions at high energy is a nonextensive process.
Keywords: $\alpha$-cluster structure; nuclear structure; multiplicity distribution of nuclear fragments; nonextensive process; the Electron-Ion Collider
@@ 1 Introduction
High-energy nuclear collisions represent a significant area of research within modern physics [2, 3, 4, 5, 6]. In these interactions, numerous particles are predominantly generated within participant regions while multiple fragments are primarily emitted from spectator regions when available. A series of recent research achievements focusing on the field of nuclear fragmentation—spanning model construction, experimental planning, and mechanism analysis—comprehensively demonstrate the vigorous development trend of this field, providing multiple perspectives for a deeper understanding of nuclear structure and nuclear reaction dynamics [7, 8, 9, 10, 11, 12, 13].
Relevant research content includes, but is not limited to, the phenomenon of nuclear fragmentation in the intermediate energy region, the influence of nuclear collective excitation modes on nuclear fragmentation, the competition between nuclear fragmentation and electromagnetic fragmentation, how nuclear micro-excited states affect the production of macroscopic nuclear fragments, and the application of Monte Carlo methods in nuclear fragmentation studies [7, 8, 9, 10, 11, 12, 13]. These studies complement and advance one another, not only propelling the field of nuclear fragmentation forward but also laying a solid foundation for progress in related areas such as nuclear physics and nuclear technology applications.
During multi-fragment emission processes, it is anticipated that spectators will form an excited nucleus. This excited nucleus subsequently undergoes fragmentation into various components [14, 15, 16, 17]. Such fragmentation reveals rich internal structures within the nucleus where protons and neutrons can combine appropriately to create intermediate configurations [18, 19, 20, 21]. When two protons and two neutrons coalesce into such an intermediate structure, it is referred to as an alpha-cluster ($\alpha$-cluster) structure [22, 23, 24, 25]. It is possible that two or more $\alpha$-cluster structures are existent in a heavy nucleus [26, 27, 28, 29, 30]. According to the $\alpha$-cluster model [31, 32, 33, 34, 35], the $\alpha$-cluster structure should have much higher probability than other intermediate structures.
@@
Furthermore, nuclear fragmentation is a non-thermal equilibrium process, rendering Boltzmann-Gibbs statistics inapplicable. Instead, one may adopt two-, three-, or multi-component distributions, where each component represents a state of local equilibrium that can be described within the framework of Boltzmann-Gibbs statistics. Consequently, a multi-temperature pattern, or temperature fluctuations, can be observed in nuclear fragmentation processes. Empirically, such multi-component distributions can be effectively fitted using Tsallis statistics, which indicates that nuclear fragmentation is inherently a nonextensive process. Therefore, nonextensive parameters derived from Tsallis statistics, including generalized temperature, entropy index, and $q$-entropy, can be employed to characterize and analyze nuclear fragmentation phenomena.
In this work, both equal and unequal probability partitioning methods are employed to derive various configurations of nuclear fragments from excited states such as 9Be, 12C, and 16O, which are expected to result from electron-nucleus ($eA$) collisions at the EIC [40]. The multiplicity distributions for all fragments as well as those with charge $Z$ will subsequently be obtained. In particular, the probability of multi-$\alpha$ or multi-He configuration or channel can be obtained, which may serve as the baseline for judging about the $\alpha$-cluster structure. In addition, in the framework of Tsallis statistics, the nonextensive generalized temperature, entropy index, and $q$-entropy are obtained from the multiplicity distribution of nuclear fragments with given charge number.
The remainder of this paper is structured as follows. Various configurations of nuclear fragments are described in Section 2. Multiplicity distributions of nuclear fragments are presented in Section 3. In Section 4, we show nonextensive parameters from multiplicity distributions of nuclear fragments. Finally, we give summary and conclusion of this work in Section 5.
@@ 2 Various configurations of nuclear fragments
$e+^{10}{\rm Be}\longrightarrow\bigg\{\begin{aligned} &(e+n)+^{9}{\rm Be}^{*}\\ &(e+p)+^{9}{\rm Li}^{*},\ \ ^{9}{\rm Li}^{*}\longrightarrow 2n+^{7}{\rm Li}^{*}\ \ {\rm or}\ \ 3n+^{6}{\rm Li}^{*},\end{aligned}$ (1)
$e+^{13}{\rm C}\longrightarrow\bigg\{\begin{aligned} &(e+n)+^{12}{\rm C}^{*}\\ &(e+p)+^{12}{\rm B}^{*},\ \ ^{12}{\rm B}^{*}\longrightarrow n+^{11}{\rm B}^{*}\ \ {\rm or}\ \ 2n+^{10}{\rm B}^{*},\end{aligned}$ (2)
@@
$e+^{17}{\rm O}\longrightarrow\bigg\{\begin{aligned} &(e+n)+^{16}{\rm O}^{*}\\ &(e+p)+^{16}{\rm N}^{*},\ \ ^{16}{\rm N}^{*}\longrightarrow n+^{15}{\rm N}^{*}\ \ {\rm or}\ \ 2n+^{14}{\rm N}^{*}.\end{aligned}$ (3)
@@
Considering the Fermi momentum of a nucleon within the nucleus, which is approximately 0.25 GeV/c/c, and the momentum per nucleon of the incident nucleus being 10 GeV/c/c, decay protons and other fragments are expected to be emitted within a forward cone characterized by a polar angle $\theta_{0}=25$ mrad. This corresponds to a pseudorapidity of $\eta=-\ln\tan(\theta_{0}/2)=4.38$. Furthermore, recoil protons can be distinguished from decay protons since recoil protons participate in multi-scattering processes which result in much larger scattering angles. It is assumed that this emission will span a wide range from nearly 0 (corresponding to $\eta=\infty$) up to approximately $10\theta_{0}$ (which corresponds to $\eta=2.07$). Indeed, it cannot be excluded that some recoil protons may have very small scattering angles, leading to exceptionally large pseudorapidities due to the influence of leading nucleons.
If the two types of protons are assumed to emit isotropically in their respective rest frames, they approximately follow Gaussian $\eta$ distributions with a common standard deviation ($\sigma_{\eta}\approx 0.91$). The decay protons predominantly distribute within the range of $4.38<\eta<4.38+4\sigma_{\eta}=8.02$, while the recoil protons primarily occupy the range of $2.07<\eta<2.07+4\sigma_{\eta}=5.71$. It is evident that there exists some overlap between decay and recoil protons in the forward cone. Although most recoil protons may have emission angles exceeding 25 mrad due to multiple scattering processes, we cannot entirely dismiss the possibility of them appearing within the forward cone. To more effectively distinguish between decay and recoil protons, one could study their energies; generally, the energy of a decay proton is nearly equal to that of an incident nucleus per nucleon, whereas the energy of a recoil proton should be lower.
@@
Table 1: The multiplicity, $N_{F}$, of all fragments and the multiplicity, $N_{Z}$, of the fragments with charge $Z$ in various configurations in excited 9Be fragmentation, where only the charge conservation is considered in the fragmentation. In the equal probability partitioning method, the fragment Be is defaulted with 50% probability to be 8Be, and in the unequal probability partitioning method, the fragment Be is defaulted with a given chance $\{M_{2}(2{\rm He})/[M_{2}(2{\rm He})+M_{2}({\rm Be})]=1/3\}$ to be 8Be, where 8Be is unstable and can decay into 2He, which is listed in the bracket, and causes $N_{F}$ to $N_{F}+1$ and $N_{Z=2}$ to $N_{Z=2}+2$. Here, the changeable $N_{F}$ and $N_{Z=2}$ are shown in the table by +1+1 and +2+2, respectively. The probabilities $f_{1}$ and $f_{2}$ of each channel obtained by the equal and unequal partitioning methods are listed respectively.
@@ 3 Multiplicity distributions of nuclear fragments
After systematically deriving distinct fragmentation channels using established partitioning methods, we conduct a fragment-specific charge analysis within each channel. This involves isolating individual fragments to precisely measure and validate their charge states, leaving no detail unexamined. Simultaneously, we meticulously track two key metrics: the total fragment multiplicity per channel and the specific multiplicities of fragments with predefined target charges. By applying the corresponding weights from the partitioning methods to the multiplicity data, we accurately determine the final multiplicity distributions for both the complete fragment set and charge-specified subsets, thereby revealing the inherent probabilistic nature of nuclear fragmentation processes.
@@
The normalized multiplicity distributions, $(1/n)(dn/dN_{F})$ or $(1/n)(dn/dN_{Z})$, in excited 9Be, 12C, and 16O fragmentation are displayed in Figures 1, 2, and 3, respectively. The solid (dashed) histograms represent the results from the equal (unequal) probability partitioning method. Figures 1(a), 2(a), and 3(a) are for the multiplicity distributions of all fragments. The multiplicity distributions of the fragments with different $Z$ are shown in different panels, where Z=1Z=1–4 in Figures 1(b)–1(e), Z=1Z=1–6 in Figures 2(b)–2(g), and Z=1Z=1–8 in Figures 3(b)–3(i) are for excited 9Be, 12C, and 16O fragmentation, respectively. It is worth noting that the predicted frequency distributions in Figures 1–3 are precise numerical values (fractions) in each bin under a given scenario.
@@
In this context, we assume that the experimental percentage of multi-He events follows a Gaussian distribution with standard deviation $\sigma$, predominantly concentrated within the range $0,4\sigma$] and centered around an expected value of $2\sigma$, which serves as our baseline. If the experimental percentage exceeds $4\sigma$, defined as twice the baseline and serving as our threshold for judgment, one can draw conclusions regarding the existence of 3He or $\alpha$ clustering with over 95% confidence level. According to this line, we conclude that 3He or $\alpha$ clustering exists in excited 9Be formed during peripheral collisions between 10B and nuclear emulsion at $E_{\rm beam}=1$ GeV/nucleon.
@@
To ensure an accurate description as possible, the heaviest fragments produced in an event—considered remnants of excited nuclear fragmentation—should be excluded from analysis. For genuine evaporation products, it must be acknowledged that they originate from the fragmentation process involving a smaller excited nucleus; thus, the partitioning methods should be reapplied specifically for this smaller nucleus. Whether a liquid-gas phase transition occurs in the overall or local area, the proportion of light fragments with $Z=1$ should exceed twice the baseline values when applying Gaussian distribution to the considered probabilities. Furthermore, if experimental measurements fall within the theoretical uncertainty range, fragmentation may be interpreted as a consequence of a general stochastic process.
Considering that excited nuclei undergo liquid-gas phase transitions locally or globally, we take the probability of the fragmentation channels including 2H–4H (2H or 4H) in 9Be, 3H–6H (3H, 4H, or 6H) in 12C, and 4H–8H (4H, 5H, 6H, or 8H) in 16O, from Tables 1–3, as the baseline values. Based on the baseline values, the fractions of channels including 2H–4H (3H–6H or 4H–8H) in excited 9Be (12C or 16O) fragmentation should be higher than 80% (54.55% or 45.45%) if the equal probability partitioning method is considered, or 58.33% (15.56% or 3.82%) if the unequal probability partitioning method is considered. Here, these percentages are derived by doubling the baseline values. The baseline values are assumed to be primarily concentrated within the range $0,4\sigma]$ when a Gaussian probability distribution with width $\sigma$ is applied.